A sample of gas with an initial volume of 28.4 L at a pressure of 735 mmHg and a temperature of 309 K is compressed to a volume of 14.5 L and warmed to a temperature of 377 K .What is the final pressure of the gas?
Answer #1

Given:

Pi = 735 mmHg

Vi = 28.4 L

Vf = 14.5 L

Ti = 309 K

Tf = 377 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(735 mmHg*28.4 L)/(309.0 K) = (Pf*14.5 L)/(377.0 K)

Pf = 1756 mmHg

= 1756/760 atm

= 2.31 atm

Answer: 2.31 atm